OpenStudy (anonymous):

compute the derivative y=e^x(x+1)^3

OpenStudy (anonymous):

use the product rule \((fg)'=f'g+g'f\) with \(f(x)=e^x,f'(x)=e^x,g(x)=(x+1)^3, g'(x)=3(x+1)^2\)

OpenStudy (anonymous):

ok thanks I didn't know what rule to use.